# Debiasing surgeon: fantastic weights and how to find them

### v1

@@ 3.4 Overview of FFW
FFW is synthesized in Alg. 1. In our approach, we hypothesize that, given a sufficiently trained and parametrized network (namely, not in an under-fitting regime), a set of spurious and target features are learned and blended together. To this end, we target to learn a mask on the encoder $\mathcal{E}$ such that the information on the bias is filtered and not usable by the task classifier $\mathcal{C}$. We therefore initialize the gating parameters $m_{i}$ to one and replace the parameters in $\mathcal{E}$ by the expression (8), where $w_{i}$ is non-trainable but the only learnable parameter is $m_{i}$. The temperature $\tau$ is initialized to one, and a bias extraction head $\mathcal{P}$ is initialized and attached to the output of $\mathcal{E}$. Then, the bias extraction head $\mathcal{P}$ and the gating parameters are trained on the training set $\mathcal{D}_{\text{train}}$ until a plateau on a validation set $\mathcal{D}_{\text{val}}$ is reached. At this point, $\tau$ is rescaled by a factor $2$ and the process is iterated until the performance of the model matches the one of the model obtained with $\tau=0$, which is our stop criterion. In the next section, we will provide our empirical findings.
@@ 4.1 Experimental setup
For our experiments, we employ architectures typically used for benchmarking bias in the three mainstream datasets Biased MNIST, CelebA, and Corrupted CIFAR10 (that will be detailed in Sec. 4.2). Specifically, for Biased MNIST, we employ the same fully convolutional network used for Rebias [5], composed of four convolutional layers with $7\!\times\!7$ kernels, with batch normalization layers. The training procedure to get the vanilla model is the same as in [53]. For the experiments on CelebA and Corrupted-CIFAR10, we used a pre-trained Resnet-18 on ImageNet, and we employed the same optimization strategy as in [46, 30]. For our FFW, we employ in all our experiments $\gamma=10$ and we use a fixed learning rate for $\mathcal{P}$ of $0.1$, until plateauing on the validation set. Noteworthily, we will employ two splits of the training set: one, biased, used for the vanilla training and indicated as $\mathcal{D}^{b}$, and another, unbiased and used by FFW to extract unbiased sub-networks, indicated as $\mathcal{D}^{u}$. In all our tables, we will report three values for our experiments: “Task” indicates the top-1 accuracy on the target task, “Bias” indicates the performance of $\mathcal{P}$ and acts as an upper bound on the information related to the bias information (as top-1 accuracy) usable by the task classifier $\mathcal{C}$, and the “Sparsity” of the extracted sub-network. Our code, provided in the Supp. Mat. will be open-sourced upon acceptance of the article.
@@ 4.2 Datasets
Below, we provide an overview of the datasets used for quantitative evaluation, chosen to represent various biases, ranging from the straightforward Biased MNIST to datasets with multiple synthetic and real-world biases. As stated in 4.1, we chose datasets that provide bias labels and will train our gating weights on balanced, unbiased datasets.
Biased MNIST. Our first dataset, introduced by Bahng et al. [5], comprises 60k samples where MNIST handwritten digits are colored with a correlation $\rho$ between color and digits. Each digit is assigned a specific color, and the samples receive a background color accordingly. We test four levels of color-digit correlation: $\rho$ = {0.99, 0.995, 0.997, and 0.999}. The bias effect, namely the background color, is assessed by testing on an unbiased dataset with $\rho$ = 0.1. As this dataset can be generated artificially by injecting a background color to black-and-white digit, for the following experiments, we built two 60k datasets for each experiment: the biased one $\mathcal{D}^{b}$ with a correlation $\rho$ and an unbiased one that constitutes $\mathcal{D}^{u}$ that we split in three parts of proportions 60%-20%-20% to obtain our $\mathcal{D}^{u}_{\text{train}},\mathcal{D}^{u}_{\text{val}},$ and $\mathcal{D}^{u}_{\text{test}}$.
CelebA. CelebA [41], a real-world dataset commonly used for debiasing evaluations, consists of 203k face images annotated with 40 attributes. Here, we focus on the classification of whether the individual has "Blond Hair" or not, where the primary bias stems from gender, with a strong tendency for "Females" to have the "Blond hair" attribute.
Corrupted CIFAR10. In the same process followed for Biased MNIST, Corrupted CIFAR10 consists in the injection of a specific correlation between an attribute (here a kind of image corruption, such as "motion blur", or "fog") and one of the ten classes of CIFAR10. The corruptions are taken from [28] and the specific corruption protocol followed here was taken from [8]. We tested their four levels of correlations: $\rho$ = {0.95, 0.98, 0.99, 0.995}.
@@ 4.3 Preliminary analysis
Table 1: Results for FFW applied on Biased-MNIST ($\rho=0.99$) with different $\gamma$ on the validation set.
$\gamma$ 2 5 10 20 50 200
Task $98.17\pm 0.27$ $98.07\pm 0.24$ $97.79\pm 0.30$ $97.52\pm 0.10$ $93.68\pm 3.77$ $23.31\pm 19.57$
Bias $21.49\pm 5.68$ $15.94\pm 1.99$ $15.74\pm 3.26$ $17.10\pm 1.93$ $16.02\pm 0.58$ $10.61\pm 1.25$
@@
Method Metric $\rho$
@@
Vanilla Task 88.46 74.23 46.04 10.02
Bias 98.45 99.75 99.96 100.00
Rubi [11] Task 93.6 43.0 90.4 13.7
EnD [56] Task 96.0 93.9 83.7 52.3
BCon+BBal [30] Task 98.1 97.7 97.3 94.0
ReBias [5] Task 88.4 75.4 65.8 26.5
LearnedMixin [14] Task 88.3 78.2 50.2 12.1
LfF [46] Task 95.1 90.3 63.7 15.3
SoftCon [30] Task 95.2 93.1 88.6 65.0
FFW Structured Task 97.63 96.68 92.46 19.91
Bias 16.89 19.03 22.59 17.57
Sparsity 45.78 47.07 38.40 47.24
FFW Task 97.79 97.30 89.36 22.29
Bias 15.74 17.14 27.84 23.13
Sparsity 46.50 44.62 36.47 43.11
We propose here a preliminary analysis where we analyze the impact of $\gamma$ on both the target task and the bias. We conduct this analysis on Biased-MNIST in the unstructured pruning setup. The results are reported in Tab. 1 and they are averaged on three seeds. As expected, we observe that for extremely large values of $\gamma$ the performance on the target task drops significantly, as the sparsity increases. However, we also notice that for values until $\gamma=20$ the task accuracy remains high. The accuracy on the bias extraction remains consistently below 20% for $\gamma>2$, which drives our selection to an intermediate value of $\gamma=10$ for all the other experiments.
4.4 Main results
In this section, we present and discuss the results obtained on the three datasets introduced in Sec. 4.2.
Tab. 2 reports the results obtained for Biased-MNIST at different proportions of bias-misaligned samples (indicated as $\rho$). For the lower values of $\rho$ we can clearly notice that both FFW structured and unstructured are among other state-of-the-art approaches, suggesting that proper pruning of vanilla-trained models, without further retraining, can yield well-generalizing performance. For very high values of $\rho$, however, we observe that FFW is falling behind other approaches. This can be explained by the fact that since the correlation between bias and target is extremely high, the disentangled information from the target is not learned by the vanilla model, due to the scarcity of unbiased representants in the training set. In Sec. 4.5 we will ablate on the need for proper fitting of all the samples in the dataset.
Table 3: Results for CelebA with different attributes.
Method Metric Correlated attributes
Gender-BlondHair Gender-Glasses
Vanilla Task 80.42 98.42
Bias 72.22 50.00
EnD [56] Task 86.9 -
LNL [35] Task 80.1 -
DI [61] Task 90.9 -
BCon+BBal [30] Task 91.4 -
Group DRO [51] Task 85.4 -
LfF [46] Task 84.2 -
FFW Task 87.08 98.58
Bias 50.00 48.66
Sparsity 49.11 46.58
FFW Structured Task 86.95 98.28
Bias 48.00 49.35
Sparsity 49.11 46.43
@@ 4.4 Main results
Method Metric Proportion of bias-misaligned samples
0.5% 1% 2% 5%
Vanilla Task $20.50\pm 0.45$ $23.90\pm 1.00$ $29.92\pm 0.10$ $43.03\pm 1.13$
Bias $89.67\pm 0.92$ $85.27\pm 1.33$ $82.18\pm 0.10$ $79.22\pm 0.45$
EnD [55] Task 19.38 23.12 34.07 36.57
HEX [60] Task 13.87 14.81 15.20 16.04
ReBias [6] Task 22.27 25.72 31.66 43.43
LfF [45] Task 28.57 33.07 39.91 50.27
DFA [37] Task 29.95 36.49 41.78 51.13
FFW Task $51.57\pm 3.16$ $58.25\pm 1.46$ $58.8\pm 0.71$ $61.48\pm 0.44$
Bias $10.03\pm 0.35$ $9.77\pm 0.06$ $10.85\pm 0.44$ $10.47\pm 0.51$
Sparsity $6.93\pm 0.60$ $2.73\pm 0.21$ $3.19\pm 0.02$ $5.32\pm 0.12$
FFW Structured Task $52.02\pm 3.24$ $57.92\pm 0.89$ $59.20\pm 1.26$ $60.97\pm 0.40$
Bias $10.65\pm 0.93$ $10.38\pm 0.43$ $10.43\pm 0.19$ $10.40\pm 0.50$
Sparsity $5.26\pm 3.39$ $2.33\pm 0.88$ $3.14\pm 0.06$ $5.31\pm 0.14$
Finally, Tab. 4 reports the results on Corrupted-CIFAR10. In this case, we have averaged our performance on three seeds, motivated by a surprising result: the performance of the vanilla model pruned through FFW surpasses state-of-the-art approaches, by a consistent margin: for 0.5% of proportion for bias-misaligned samples, FFW beats other techniques by approximately 20%. We hypothesize that the corruptions present in this dataset are easy to remove by pruning; still, it is surprising that the vanilla model, properly pruned and without extra fine-tuning, can reach such performance.
@@ 4.5 Importance of fitting the training set
Table 5: Results for FFW Biased-MNIST with $\rho=0.99$ with different extraction times. Note that one full training consists of 600 batches (therefore the three first columns represent respectively 10, 100, and 300 iterations).
Accuracy Epochs
@@
As an ablation study, we propose to extract unbiased subnetworks from a vanilla-trained model from models trained for a different amount of epochs. For this study, we select Biased-MNIST with $\rho=0.99$. We report the results in Tab. 5, averaging the results on three seeds. For this setup, we evidently observe that, even training less than an epoch (as little as 1/60 of an epoch) the vanilla model perfectly fits the bias, and the disentangled features are completely not learned, until the completion of the first epoch. Progressively though, non-biased features are learned by the model, and the accuracy of the target task increases. This suggests that a simple approach as FFW can work on vanilla-trained models if they are sufficiently trained (and fitting) on the target task.
@@ 5 Conclusion
Secondly, we have proposed FFW, a technique that, without requiring fine-tuning the vanilla model, is able to surgically remove parameters from the model, unveiling the existence of an unbiased sub-network. We have proposed both an unstructured and a structured variant of such an approach, which also provides guarantees on the biased information employable for solving the target task. On three common benchmarks, we have observed that such sub-networks exist, leading to performance comparable with other state-of-the-art approaches. One major finding is that they are even structured, potentially leading also to computational gains. This finding bridges the sparsity and debiasing communities, opening the road to the design of more energy-efficient debiasing approaches.
@@ Acknowledgements
Part of this work was funded by the French National Research Agency (ANR-22-PEFT-0007) as part of France 2030 and the NF-FITNESS project, and by Hi!PARIS Center on Data Analytics and Artificial Intelligence. Besides, this work was also funded by the European Union’s Horizon Europe research and innovation program under grant agreement No. 101120237 (ELIAS).
Appendix A More experiments
A.1 Testing our method on multiple biases: Multi-Color MNIST
Table 6: Test accuracy on four subsets of Multi-Color MNIST. The “Unbiased” one is the average of the four. A stands for Aligned and C for Conflicting (regarding the bias-alignement of the samples in each subset).
Method Sparsity Bias acc. [%] (normal-$\downarrow$) Task acc. [%] (normal-$\uparrow$)
Vanilla 0.00 96.38 64.7 100.0 97.7 30.37 7.68 58.9
LfF [46] 0.00 - - 99.6 4.7 98.6 5.1 52.0
EIIL [17] 0.00 - - 100.0 97.2 70.8 10.9 69.7
PGI [1] 0.00 - - 98.6 82.6 26.6 9.5 54.3
DebiAN [40] 0.00 - - 100.0 95.6 76.5 16.0 72.0
VCBA [44] 0.00 - - 100.0 90.9 77.5 24.1 $\boldsymbol{73.1}$
FFW 16.95 20.27 16.55 34.57 35.17 39.86 $\boldsymbol{35.85}$ 36.37
To complexify the challenge posed by Biased MNIST, Li et al. proposed a bi-colored version in [40] to assess models’ performance on multiple biases simultaneously. In this version, the left and right sides of the background have different colors, each correlated to the target with $\rho_{L}$ and $\rho_{R}$ respectively. We adopt their setup with $\rho_{L}$ = 0.99 and $\rho_{R}$ = 0.95. We applied FFW to this dataset to obtain the results presented in Tab. 6. If FFW yields worst results on average than the vanilla network, we can notice that it is the pruning led to the best results on the worst sub-group: the one with bias-conflicting backgrounds on both sides. On this one, we get an increase of 11%. We can also notice that the results are comparable on every subset, meaning that the network is nos its choices on the background color anymore.
Appendix B Visualizations
B.1 Is FFW helping the network focus on the right features ?
Figure 3: Grad-Cam visualization of the effects of FFW on Biased-MNIST with $\rho=0.997$.

### v2

@@ 3.4 Overview of FFW
FFW is synthesized in Alg. 1. In our approach, we hypothesize that, given a sufficiently trained and parametrized network (namely, not in an under-fitting regime), a set of spurious and target features are learned and blended together. To this end, we target to learn a mask on the encoder $\mathcal{E}$ such that the information on the bias is filtered and not usable by the task classifier $\mathcal{C}$. We therefore initialize the gating parameters $m_{i}$ to $1$ and replace the parameters in $\mathcal{E}$ by the expression (8), where $w_{i}$ is non-trainable but the only learnable parameter is $m_{i}$. The temperature $\tau$ is initialized to $1$, and a bias extraction head $\mathcal{P}$ is initialized and attached to the output of $\mathcal{E}$. Then, $\mathcal{P}$ and the gating parameters are trained on the training set $\mathcal{D}_{\text{train}}$ until a plateau on a validation set $\mathcal{D}_{\text{val}}$ is reached. At this point, $\tau$ is rescaled by a factor $2$ and the process is iterated until the performance of the model matches the one of the model obtained with $\tau=0$, which is our stop criterion.We will provide our empirical findings below.
@@ 4.1 Experimental Setup
For our experiments, we employ architectures typically used for benchmarking bias in the four datasets Biased MNIST, CelebA, Corrupted CIFAR10 and Multi-Color MNIST (described in Sec. 4.2). Specifically, for Biased MNIST, we employ the same fully convolutional network used for [4], composed of four convolutional layers with $7\!\times\!7$ kernels, with batch normalization layers. The training procedure to get the vanilla model is the same as in [50]. For the experiments on CelebA and Corrupted-CIFAR10, we used a pre-trained Resnet-18 on ImageNet, and we employed the same optimization strategy as in [43, 27]. For our FFW, we employ in all our experiments $\gamma=10$ and we use a fixed learning rate for $\mathcal{P}$ of $0.1$, until plateauing on the validation set. Noteworthily, we will employ two splits of the training set: one, biased, used for the vanilla training and indicated as $\mathcal{D}^{b}$, and another, unbiased and used by FFW to extract unbiased sub-networks, indicated as $\mathcal{D}^{u}$. In all our tables, we report three values for our experiments: “Task” indicates the top-1 accuracy on the target task, “Bias” indicates the performance of $\mathcal{P}$ and acts as an upper bound on the information related to the bias information (as top-1 accuracy) usable by $\mathcal{C}$, and the sparsity of the extracted sub-network (sometimes displayed as “Spars.”). Our code is accessible at this link.
@@ 4.2 Datasets
Below, we provide an overview of the datasets used for quantitative evaluation, chosen to represent various biases types (synthetic or real-world) and number. As stated in 4.1, we chose datasets that provide bias labels and train our gating weights on balanced, unbiased datasets.
Biased MNIST. Introduced by Bahng et al. [4], comprises 60k samples where MNIST handwritten digits are colored with a correlation $\rho$ between color and digits. Each digit is assigned a specific color, and the samples receive a background color accordingly. We test four levels of color-digit correlation: $\rho$ = {0.99, 0.995, 0.997, and 0.999}. The bias effect, namely the background color, is assessed by testing on an unbiased dataset: $\rho$ = 0.1. As this dataset can be generated artificially by injecting a background color to black-and-white digit, for the following experiments, we built two 60k datasets for each experiment: the biased $\mathcal{D}^{b}$ with a correlation $\rho$ and an unbiased $\mathcal{D}^{u}$ that we split in three parts of proportions 60%-20%-20% to obtain our $\mathcal{D}^{u}_{\text{train}},\mathcal{D}^{u}_{\text{val}},$ and $\mathcal{D}^{u}_{\text{test}}$.
CelebA. CelebA [38], a real-world dataset commonly used for debiasing evaluations, consists of 203k face images annotated with 40 attributes. Here, we focus on the classification of whether the individual has "BlondHair" or not, where the primary bias stems from the strong tendency for perceived females to be blond.
Corrupted CIFAR10. Similarly to Biased MNIST, Corrupted CIFAR10 consists in the injection of a specific correlation between an attribute (here an image corruption, such as "motion blur", or "fog") and one of the ten classes of CIFAR10. The corruptions are taken from [25] and the corruption protocol was taken from [6]. We tested their levels of correlation: $\rho$ = {0.95, 0.98, 0.99, 0.995}.
Multi-Color MNIST. To complexify the challenge posed by Biased MNIST, Li et al. proposed a bi-colored version in [37] to assess models’ performance on multiple biases simultaneously. In this version, the left and right sides of the background have different colors, each correlated to the target with $\rho_{L}$ and $\rho_{R}$ respectively. We adopt their setup with $\rho_{L}$ = 0.99 and $\rho_{R}$ = 0.95.
@@ 4.3 Preliminary analysis
Table 1: Results for FFW applied on Biased-MNIST ($\rho=0.99$) for different $\gamma$.
Metric Regularization coefficient $\gamma$
2 5 10 20 50 200
Task Accuracy $98.17\pm 0.27$ $98.07\pm 0.24$ $97.79\pm 0.30$ $97.52\pm 0.10$ $93.68\pm 3.77$ $23.31\pm 19.57$
Bias Accuracy $21.49\pm 5.68$ $15.94\pm 1.99$ $15.74\pm 3.26$ $17.10\pm 1.93$ $16.02\pm 0.58$ $10.61\pm 1.25$
@@
We propose here a preliminary analysis where we analyze the impact of $\gamma$ on both the target task and the bias. We conduct this analysis on Biased-MNIST in the unstructured pruning setup. The results are reported in Tab. 1 and they are averaged on three seeds. As expected, we observe that for extremely large values of $\gamma$ the performance on the target task drops significantly, as the sparsity increases. However, we also notice that for values until $\gamma=20$ the task accuracy remains high. The bias accuracy remains consistently below 20% for $\gamma>2$, which drives our selection to an intermediate value of $\gamma=10$ for the other experiments.
4.4 Main results
In this section, we discuss our results on the datasets introduced in Sec. 4.2.
@@ 4.4 Main results
Method Proportion of bias-aligned samples: $\rho$
@@
Task Bias Spars. Task Bias Spars. Task Bias Spars. Task Bias Spars.
Vanilla 88.46 98.45 - 74.23 99.75 - 46.04 99.96 - 10.02 100.0 -
Rubi [9] 93.6 - - 43.0 - - 90.4 - - 13.7 - -
EnD [52] 96.0 - - 93.9 - - 83.7 - - 52.3 - -
BCon+BBal [27] 98.1 - - 97.7 - - 97.3 - - 94.0 - -
ReBias [4] 88.4 - - 75.4 - - 65.8 - - 26.5 - -
LearnedMixin [12] 88.3 - - 78.2 - - 50.2 - - 12.1 - -
LfF [43] 95.1 - - 90.3 - - 63.7 - - 15.3 - -
SoftCon [27] 95.2 - - 93.1 - - 88.6 - - 65.0 - -
FFW Structured 97.63 16.89 45.78 96.68 19.03 47.07 92.46 22.59 38.40 19.91 17.57 47.24
FFW 97.79 15.74 46.50 97.30 17.14 44.62 89.36 27.84 36.47 22.29 23.13 43.11
Tab. 2 reports the results obtained for Biased-MNIST at different proportions of bias-misaligned samples (indicated as $\rho$). For the lower values of $\rho$ we can clearly notice that both FFW structured and unstructured are among other state-of-the-art approaches, suggesting that proper pruning of vanilla-trained models, without further retraining, can yield well-generalizing performance. For $\rho=0.999$, however, we observe that FFW is falling behind other approaches. This can be explained by the fact that the scarcity of unbiased representants in the training set prevents the vanilla model from learning the bias-disentangled information. In Sec. 4.5 we display the need to properly fit the training data.
Table 3: Results for FFW applied on CelebA with different target attributes.
Method Target task attribute
BlondHair Glasses
Task Bias Sparsity Task Bias Sparsity
Vanilla 80.42 72.22 0.00 98.42 50.00 0.00
EnD [52] 86.9 - - - - -
LNL [33] 80.1 - - - - -
DI [56] 90.9 - - - - -
BCon+BBal [27] 91.4 - - - - -
Group DRO [48] 85.4 - - - - -
LfF [43] 84.2 - - - - -
LWBC [34] 85.5 - - - - -
FFW Structured 86.95 48.00 49.11 98.28 49.35 46.43
FFW 87.08 50.00 49.11 98.58 48.66 46.58
@@
Method Proportion of bias-aligned samples: $\rho$
0.995 0.99 0.98 0.95
Task Bias Spars. Task Bias Spars. Task Bias Spars. Task Bias Spars.
Vanilla 20.50 89.67 0.00 23.90 85.27 0.00 29.92 82.18 0.00 43.03 79.22 0.00
EnD [52] 19.38 - - 23.12 - - 34.07 - - 36.57 - -
HEX [55] 13.87 - - 14.81 - - 15.20 - - 16.04 - -
ReBias [4] 22.27 - - 25.72 - - 31.66 - - 43.43 - -
LfF [42] 28.57 - - 33.07 - - 39.91 - - 50.27 - -
DFA [35] 29.95 - - 36.49 - - 41.78 - - 51.13 - -
uLA [53] 34.39 - - 62.49 - - 63.88 - - 74.49 - -
FFW 51.57 10.03 6.93 58.25 9.77 2.73 58.80 10.85 3.19 61.48 10.47 5.32
FFW Structured 52.02 10.65 5.26 57.92 10.38 2.33 59.20 10.43 3.14 60.97 10.40 5.31
Tab. 4 reports the results on Corrupted-CIFAR10. In this case, the vanilla model pruned through FFW surpasses state-of-the-art approaches, by a high margin in some cases: for $\rho=0.05$, FFW beats other techniques by approximately 17%. We hypothesize that the corruptions present in this dataset are easy to remove by pruning; still, it is surprising that the vanilla model, properly pruned and without extra fine-tuning, can reach such performance.
Table 5: Results for FFW applied on Multi-Color MNIST. The accuracies are computed on four subsets of the dataset. The “Unbiased” one is the average of the four. $A$ and $C$ stand for Aligned and Conflicting (regarding the samples).
Method Sparsity [%] ($\uparrow$) Bias acc. [%] ($\downarrow$) Task acc. [%] ($\uparrow$)
Vanilla 0.00 96.38 64.7 100.0 97.7 30.37 7.68 58.9
LfF [43] - - - 99.6 4.7 98.6 5.1 52.0
EIIL [14] - - - 100.0 97.2 70.8 10.9 69.7
PGI [2] - - - 98.6 82.6 26.6 9.5 54.3
DebiAN [37] - - - 100.0 95.6 76.5 16.0 72.0
VCBA [41] - - - 100.0 90.9 77.5 24.1 $\boldsymbol{73.1}$
FFW 16.95 20.27 16.55 34.57 35.17 39.86 $\boldsymbol{35.85}$ 36.37
Finally, Tab. 5 displays our results on Multi-Color MNIST. If FFW yields worst results on average than the vanilla network, we can notice that it led to the best results on the worst sub-group: the one with bias-conflicting backgrounds on both sides: we get an increase of 11%. We can also notice that the results are comparable on every subset, meaning that the network doesn’t base its choices on the background color anymore.
@@ 4.5 Importance of fitting the training set
Table 6: Results for Biased-MNIST with $\rho=0.99$ with different extraction times. Note that one full training epoch consists of 600 batches (therefore the three first columns represent respectively 10, 100, and 300 iterations).
Method Metric Epochs
@@
As an ablation study, we propose to extract unbiased sub-networks from a vanilla-trained model from models trained for a different amount of epochs. For this study, we select Biased-MNIST with $\rho=0.99$. We report the results in Tab. 6, averaging the results on three seeds. For this setup, we evidently observe that, even training less than an epoch (as little as 1/60 of an epoch) the vanilla model perfectly fits the bias, and the disentangled features are completely not learned, until the completion of the first epoch. Progressively though, non-biased features are learned by the model, and the accuracy of the target task increases. This suggests that a simple approach as FFW can work on vanilla-trained models if they are sufficiently trained (and fitting) on the target task.
@@ 5 Conclusion
Secondly, we have proposed FFW, a technique that, without requiring fine-tuning the vanilla model, is able to surgically remove parameters from the model, unveiling the existence of an unbiased sub-network. We have proposed both an unstructured and a structured variant of such an approach, which also provides guarantees on the biased information employable for solving the target task. On four common benchmarks, we have observed that such sub-networks exist, leading to performance comparable with other state-of-the-art approaches. One major finding is that they are even structured, potentially leading also to computational gains. This finding bridges the sparsity and debiasing communities, opening the road to the design of more energy-efficient debiasing approaches.
@@ Acknowledgements
Part of this work was funded by the French National Research Agency (ANR-22-PEFT-0007) as part of France 2030 and the NF-FITNESS project and by Hi!PARIS Center on Data Analytics and Artificial Intelligence. Besides, this work was also funded by the European Union’s Horizon Europe research and innovation program under grant agreement No. 101120237 (ELIAS).
Appendices
Appendix 0.A Visualizations
0.A.1 Is FFW helping the network focus on the right features ?
Figure 3: Grad-Cam visualization of the effects of FFW on Biased-MNIST with $\rho=0.997$